OpenStudy (anonymous):

Factor completely by factoring out any common factors and then factoring by grouping. 6x2 – 5xy + 6x – 5y

OpenStudy (mathstudent55):

1. Is there a common factor in all 4 terms?

OpenStudy (anonymous):

not that I see at all. shouldn't I just continue to factor?

OpenStudy (mathstudent55):

Good. 2. Is there a common factor you can take out of the first two terms? Is there a factor you can factor out of the last two terms?

OpenStudy (mathstudent55):

Can you factor a common factor out of \(6x^2 - 5xy\) Can you factor a common factor out of \(6x - 5y\)

OpenStudy (anonymous):

I can take x out of the first but nothing from the second set of terms

OpenStudy (mathstudent55):

You are correct, but here's the trick. In factoring a 4-term polynomial by grouping, you always need to factor something out even if that thing is simply 1. \(6x^2 - 5xy + 6x - 5y\) \(= x(6x - 5y) + 1(6x - 5y) \) Ok?

OpenStudy (anonymous):

ooooooooh ok, thank you

OpenStudy (mathstudent55):

3. Now you do the final step. Factor out the common factor of 6x - 5. \(= x(6x - 5y) + 1(6x - 5y)\) \( (6x - 5y)(x + 1) \)